is 2/3y+6x-3=4 equal to y=9x+10 1/2? How do you find it out?

Answer by checkley77(12844)

Solve the first equation for y & compare the results.
2/3y+6x-3=4
2Y/3=-6X+4+3
2Y/3=-6X+7
2Y=3(-6X+7)
2Y=-18X+21
Y=-18X/2+21/2
Y=-9X+10 1/2 |
y=9x+10 1/2? | THESE 2 EQUATIONS ARE NOT EQUAL.